-7x+y=-19
-2x+3y=-19
-7x-8y=9
-4x+9y=-22
5x+4y=-30
3x-9y=-18
-14=-20y-7x
10y+4=2x
3+2x-y=0
-3-7y=10x
Solve by elimination

Answer by solver91311(24713)


  -7x +  y = -19
  -2x + 3y = -19

Multiply the first equation by -3

  21x - 3y =  57
 - 2x + 3y = -19

Add the two equations

  19x + 0y =  38

Solve for x

    x      =  2

Substitute the now known value of x into either original equation

-2(2) + 3y = -19
        3y = -15
         y = - 5

Check your work:

-2(2) + 3(-5) =  -4 - 15 = -19 -- Checks
-7(2) +   -5  = -14 -  5 = -19 -- Checks

If you are wondering what happened to the other problems you posted, go back and read the instructions for posting again.


John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism